If 4g of a gas occupies a volume of 2L at STP;
(i). What is its density? (D = m/V)
(ii). What is its relative molecular mass?
(iii). Which of these gases is it most likely to be?

Cl2 CO2 NO CH4
1
Expert's answer
2020-09-24T11:22:41-0400

find the density of gas d=m/V=4/2=2 g/l

find Mr=d *Vm=2*22.4=44.8

the most likely to be CO2, because Mr(CO2)=44